### v4

# Smallest Suffixient Sets: Effectiveness, Resilience, and Calculation

Abstract
A suffixient set is a novel combinatorial object that captures the essential information of repetitive strings in a way that, provided with a random access mechanism, supports various forms of pattern matching. In this paper, we study the size $\chi$ of the smallest suffixient set as a repetitiveness measure: we place it between known measures and study its sensitivity to various string operations. As a corollary of our results, we give simple online algorithms to compute smallest suffixient sets.
1 Introduction
The study of repetitive string collections has recently attracted considerable interest from the stringology community, triggered by practical challenges such as representing huge collections of similar strings in a way that they can be searched and mined directly in highly compressed form [25, 26]. An example is the European ’1+ Million Genomes’ Initiative [^1] https://digital-strategy.ec.europa.eu/en/policies/1-million-genomes, which aims at sequencing over a million human genomes: while this data requires around 750TB of storage in raw form (using 2 bits per base), the high similarity between human genomes would allow storing it in querieable form using two orders of magnitude less space.
An important aspect of this research is to understand how to measure repetitiveness, especially when those measures reflect the size of compressed representations that offer different access and search functionalities on the collection. Various repetitiveness measures have been proposed, from abstract lower bounds to those related to specific text compressors and indices; a relatively up-to-date survey is maintained [27]. Understanding how those measures relate to each other sheds light on what search functionality is obtained at what space cost.
A relevant measure proposed recently is the size $\chi$ of the smallest suffixient set of the text collection [6], whose precise definition will be given later. Within $O(\chi)$ size, plus a random-access mechanism on the string, it is possible to support some text search functionalities, such as finding one occurrence of a pattern, or finding its maximal exact matches (MEMs), which is of central use on various bioinformatic applications [4].
While there has been some work already on how to build minimal suffixient sets and how to index and search a string within their size, less is known about that size, $\chi$, as a measure of repetitiveness. It is only known [6] that $\gamma=O(\chi)$ and $\chi=O(\overline{r})$ on every string family, where $\gamma$ is the size of the smallest string attractor of the collection (a measure that lower bounds most repetitiveness measures) [18] and $\overline{r}$ is the number of equal-letter runs of the Burrows-Wheeler Transform (BWT) [3] of the reversed string.
In this paper we better characterize $\chi$ as a repetitiveness measure. First, we study how it behaves when the string undergoes updates, showing in particular that it grows by $O(1)$ when appending or prepending symbols, but that it can grow additively by $\Omega(\log n)$ upon arbitrary edit operations or rotations, and by $\Omega(\sqrt{n})$ when reversing the string. For this last operation, we show that $\chi$ cannot grow by a factor larger than $2$, however. Second, we show that $\chi=O(r)$ on every string family, where $r$ is the number of equal-letter runs of the BWT of the string. We also show that there are string families where $\chi=o(v)$, where $v$ is the size of the smallest lexicographic parse [28] (an alternative to the size of the Lempel-Ziv parse [20], which behaves similarly). In particular, this holds on the Fibonacci strings, where we fully characterize the only 2 smallest suffixient sets of size 4, and further prove that $\chi\leq\sigma+2$ on all substrings of episturmian words over an alphabet of size $\sigma$. Since $v=O(r)$ on all string families, this settles $\chi$ as a strictly smaller measure than $r$, which is a more natural characterization than in terms of the reverse string. We also show that $\chi$ is incomparable with most “copy-paste” based measures [25], as there are families where it is strictly smaller and others where it is strictly larger than any of those measures.
This result relates to the important question of whether a measure $\mu$ is reachable (i.e., one can represent the string within $O(\mu)$ space), accessible (i.e., one can access any string position from an $O(\mu)$-size representation, in sublinear time), or searchable (i.e., one can search for patterns in sublinear time within space $O(\mu)$). Measure $r$ is, curiously, the only one to date being reachable and searchable, but not known to be accessible. Now $\chi$ emerges as a measure smaller than $r$, which can search if provided with a mechanism to efficiently access substrings ($r$ does not need access to support searches). Unlike $r$, $\chi$ is yet not known to be reachable (as its relation to the smallest known reachable measure, the size $b$ of the smallest bidirectional macro scheme [32], remains unknown). As said, it is known that $\gamma=O(\chi)$, but it is unknown whether $\gamma$ is reachable or not.
Our final contribution are new, extremely simple, online algorithms to compute smallest suffixient sets (and thus $\chi$), scanning the text left to right or right to left. While there already exist efficient algorithms to do this [5], our new algorithm can, at any point, exhibit a smallest suffixient set for the string it has just consumed. Just as online suffix tree constructions [34, 33], our algorithm uses $O(n)$ space and worst-case time in the transdichotomous RAM model over polynomial-size integer alphabets. The best previous algorithm obtains the same result [5], but it is not online and starts from the suffix array and other components of the suffix tree, whereas ours starts from the text and builds the suffix tree at the same time. All linear-time suffix tree construction algorithms run under the same model of computation.
2 Preliminaries
An ordered alphabet $\Sigma=\{a_{1},\dots,a_{\sigma}\}$ is a finite set of symbols equipped with a total order << such that $a_{1}<a_{2}<\dots<a_{\sigma}$. When $\sigma=2$, we assume $\Sigma=\{\texttt{a},\texttt{b}\}$ with $\texttt{a}<\texttt{b}$. The special symbol \$, if it appears, is always assumed to be the smallest of the alphabet.
A string $w[1\mathinner{.\,.}n]$ (or simply $w$ if it is clear from the context) of length $|w|=n$ over the alphabet $\Sigma$ is a sequence $w[1]w[2]\cdots w[n]$ of symbols where $w[i]\in\Sigma$ for all $i\in[1,n]$. The empty string of length 0 is denoted $\epsilon$. We denote by $\Sigma^{*}$ the set of all strings over $\Sigma$. Additionally, we let $\Sigma^{+}=\Sigma^{*}\setminus\{\epsilon\}$ and $\Sigma^{k}=\{w\in\Sigma^{*}\mid|w|=k\}$. We denote by $w[i\mathinner{.\,.}j]$ the substring $w[i]w[i+1]\cdots w[j]$. If $x=x[1\mathinner{.\,.}n]$ and $y=y[1\mathinner{.\,.}m]$ are strings, we define the concatenation operation applied on $x$ and $y$, as the string obtained by juxtaposing these two strings, that is, $x\cdot y=x[1]x[2]\cdots x[n]y[1]\cdots y[m]=xy$. A string $x$ is a substring of $w$ if $w=yxz$ for some $y,z\in\Sigma^{*}$. A string $x$ is a prefix of $w$ if $w=xy$ for some $y\in\Sigma^{*}$. Analogously, $x$ is a suffix of $w$ if $w=yx$ for some $y\in\Sigma^{*}$. We say that substrings, prefixes, and suffixes are non-trivial if they are different from $w$ and $\epsilon$. The set of substrings of $w$ is denoted by $\mathcal{F}_{w}$. We also let $\mathcal{F}_{w}(k)=\mathcal{F}_{w}\cap\Sigma^{k}$. The reverse of a finite string $w$ is the string $w^{R}=w[n]\cdot w[n-1]\cdots w[1]$. We denote by $\mathcal{R}(w)$ the multiset of rotations of $w[1\mathinner{.\,.}n]$, that is, $\mathcal{R}(w)=\{w[i+1\mathinner{.\,.}n]w[1\mathinner{.\,.}i]\mid i\in[1\mathinner{.\,.}n]\}$. The Burrows-Wheeler transform (BWT) of a string $w$, denoted $\mathtt{BWT}(w)$, is the transformation of $w$ obtained by collecting the last symbol of all rotations in $\mathcal{R}(w)$ in lexicographic order. The BWT matrix $B(w)$ of $w$ is the $(n\times n)$-matrix where the $i$-th row is the $i$-th rotation of $w$ in lexicographic order.
A right-infinite string w —we use boldface to emphasize its infinite length— over $\Sigma$ is any infinite sequence $\mathbb{Z}^{+}\rightarrow\Sigma$. The set of all infinite strings over $\Sigma$ is denoted $\Sigma^{\omega}$. A substring of w is the finite string $\textbf{w}[i\mathinner{.\,.}j]$ for any $1\leq i\leq j$. A prefix of w is a finite substring of the form $\textbf{w}[1\mathinner{.\,.}n]$ for some $n\geq 0$. The substring complexity function $P_{\textbf{w}}(k):\mathbb{Z}^{+}\cup\{0\}\rightarrow\mathbb{Z}^{+}$ counts the number of distinct substrings of length $k$ in w, for any $k\in\mathbb{Z}^{+}\cup\{0\}$, that is, $P_{\textbf{w}}(k)=|\mathcal{F}_{\textbf{w}}(k)|$. For a finite string $w[1\mathinner{.\,.}n]$, the domain of $P_{w}$ is restricted to $0\mathinner{.\,.}n]$.
2.1 Measures of repetitiveness
Figure 1: Relations between relevant repetitiveness measures and how our results place $\chi$ among them. An arrow $\mu_{1}\to\mu_{2}$ means that $\mu_{1}=O(\mu_{2})$ for all strings and, save for $c\to z$, $z_{no}\to z_{end}$, and $z_{end}\to z_{e}$, there is a string family where $\mu_{1}=o(\mu_{2})$. The dotted arrows mark only this last condition, so they are not transitive. Measures in light gray nodes are known to be reachable; those in dark gray are accessible and searchable; and $r$ is hatched because it is searchable but not known to be accessible.
In this work, we will relate $\chi$, in asymptotic terms, with several well-established measures of repetitiveness [25, 27]: $\delta=\max_{k\in[0\mathinner{.\,.}n]}(\mathcal{F}_{w}(k)/k)$ (a measure of substring complexity), $\gamma$ (the smallest string attractor), $b$ (the size of the smallest bidirectional macro scheme), $z$ (the size of the Lempel-Ziv parse), $z_{no}$ (the same without allowing phrases to overlap their sources), $z_{e}$ (the size of the greedy LZ-End parse), $z_{end}$ (the size of the minimal LZ-End parse), $v$ (the size of the smallest lexicographic parse), $r$ (the number of equal-letter runs in the BWT of the string), $g$ (the size of the smallest context-free grammar generating only the string), $g_{rl}$ (the same allowing run-length rules), and $c$ (the size of the smallest collage system generating only the string). Except for $\delta$, $\gamma$ and $r$, these measures are said to be copy-paste because they refer to a way of cutting the sequence into chunks that can be copied from elsewhere in the same sequence. Indeed, $\delta$ and $\gamma$ are lower-bound measures, the former known to be unreachable and the latter not known to date to be reachable; all the others are. The smallest measures known to be accessible (and searchable) are $z_{end}$ and $g_{rl}$, and $r$ is searchable but not known to be accessible.
The known relations between those measures are summarized in Fig. 1, where we have added the results we obtain in this paper with respect to $\chi$.
2.2 Edit operations and sensitivity functions
The so-called edit operations are insertion, substitution and deletion of a single character on a string. We denote $\mathtt{ins}_{\Sigma}(w)$, $\mathtt{sub}_{\Sigma}(w)$, $\mathtt{del}_{\Sigma}(w)$ the sets of strings that can be obtained by applying an edit operation to $w$. In addition, we let $\mathtt{prepend}_{\Sigma}(w)$ and $\mathtt{append}_{\Sigma}(w)$ be $\mathtt{ins}_{\Sigma}(w)$ restricted to the insertion being made at the beginning and the end of the string, respectively.
A repetitiveness measure $\mu$ is monotone or non-decreasing to the insertion of a single character if $\mu(w^{\prime})-\mu(w)\geq 0$ for any $w$ and $w^{\prime}\in\mathtt{ins}_{\Sigma}(w)$. More generally, the additive sensitivity and multiplicative sensitivity functions of a repetitiveness measure $\mu$ to the insertion of a single character are the maximum possible values of $\mu(w^{\prime})-\mu(w)$ and $\mu(w^{\prime})/\mu(w)$, respectively. We define the concept of monotonicity and sensitivity functions for the remaining string operations analogously.
3 Suffixient Sets and the Measure $\chi$
In this section we define the central combinatorial objects and measures we analyse on this work. Note that some of our definitions are slightly different from their original formulation [4, 5], because we do not always assume that all strings are \$-terminated.
Definition 1(Right-maximal Substrings and Right-extensions [4, 5])
Let $w\in\Sigma^{*}$. A substring $x$ of $w$ is right-maximal if there exist at least two distinct symbols $a,b\in\Sigma$ such that both $xa$ and $xb$ are substrings of $w$. For any right-maximal substring $x$ of $w$, the substrings $xa$ with $a\in\Sigma$ are called right-extensions. We denote the set of right-extensions in $w$ by $E_{r}(w)=\{xa\mid\exists b:b\neq a,xa\in\mathcal{F}_{w},xb\in\mathcal{F}_{w}\}$.
We distinguish a special class of right-extensions that are not suffixes of any other right-extension.
Definition 2(Super-maximal Extensions [4, 5])
The set of super-maximal extensions of $w$ is $\mathcal{S}_{r}(w)=\{x\in E_{r}(w)\mid\forall y\in E_{r}(w),y=zx\Rightarrow z=\varepsilon\}$. Moreover, we let $\mathtt{sre}(w)=|\mathcal{S}_{r}(w)|$.
We now define suffixient sets for strings not necessarily \$-terminated; we introduce later the special terminator \$.
Definition 3(Suffixient Set [4, 5])
Let $w[1\mathinner{.\,.}n]\in\Sigma^{*}$. A set $S\subseteq[1\mathinner{.\,.}n]$ is a suffixient set for $w$ if for every right-extension $x\in E_{r}(w)$ there exists $j\in S$ such that $x$ is a suffix of $w[1\mathinner{.\,.}j]$.
Intuitively, a suffixient set is a collection of positions of $1\mathinner{.\,.}|w|]$ capturing all the right-extensions appearing in $w$. The smallest suffixient sets, which are suffixient sets of minimum size, have also been characterized in terms of super-maximal right-extensions. The next definition simplifies the original one [4, 5].
Definition 4(Smallest Suffixient Set)
Let $w[1\mathinner{.\,.}n]\in\Sigma^{*}$. A suffixient set $S\subseteq[1\mathinner{.\,.}n]$ is a smallest suffixient set for $w$ if there is a bijection $pos:\mathcal{S}_{r}(w)\to S$ such that every $x\in\mathcal{S}_{r}(w)$ is a suffix of $w[1\mathinner{.\,.}pos(x)]$.
In its original formulation, the measure $\chi$ is defined over \$-terminated strings. Here, we define $\chi(w)$ with the \$ being implicit, not being part of $w$.
Definition 5(Measure $\chi$ [4, 5])
Let $w\in\Sigma^{*}$ and assume $\text{\tt\textdollar}\not\in\mathcal{F}_{w}$. Then, $\chi(w)=|\mathcal{S}|$, where $\mathcal{S}$ is a smallest suffixient set for $w\text{\tt\textdollar}$.
[…]
4 Sensitivity of $\chi$ to String Operations
The sensitivity to string operations has been studied for many repetitiveness measures [1, 9, 10, 14, 15, 24, 29, 30]. A robust repetitiveness measure should not change much upon small changes in the sequence. For instance, $b$, $z$, and $g$ can increase only by a multiplicative constant after an edit operation [1], and they can increase only by a constant additive factor when prepending or appending a character. On the other hand, $r$ can increase by a $\Theta(\log n)$ factor when appending a character [15, Prop. 37]. Other results have been obtained concerning more complex string operations, like reversing a string [14], or applying a string morphism [9, 10].
In this section we study how $\mathtt{sre}$ and $\chi$ behave in this respect. We leave for the next section our results on the multiplicative sensitivity of $\chi$ to reversing the string.
4.1 Appending and prepending symbols
We first show that $\mathtt{sre}(w)$ and $\chi(w)$ increase or decrease by only additive constants when adding or removing symbols at the extremes of $w$. We start by proving the following useful lemma.
Lemma 1
If $E_{r}(w_{1})\subseteq E_{r}(w_{2})$, then $\mathtt{sre}(w_{1})\leq\mathtt{sre}(w_{2})$.
Proof
Let $x,y\in\mathcal{S}_{r}(w_{1})$ with $x\neq y$. Because $x\in E_{r}(w_{2})$, there exists $z\in\mathcal{S}_{r}(w_{2})$ with $x$ a suffix of $z$. Because $y$ is not a suffix of $x$ and vice versa, $y$ cannot be a suffix of $z$. Therefore, the map $x\mapsto z$ with $x\in\mathcal{S}_{r}(w_{1})$, $z\in\mathcal{S}_{r}(w_{2})$, and $z=z^{\prime}x$ for some $z^{\prime}\in\Sigma^{*}$ is injective and then $\mathtt{sre}(w_{1})\leq\mathtt{sre}(w_{2})$. ∎
We now prove that $\mathtt{sre}(w)$ grows only by $O(1)$ when prepending or appending characters.
Lemma 2
Let $w\in\Sigma^{*}$, and $c\in\Sigma$. It holds $\mathtt{sre}(w)\leq\mathtt{sre}(wc)\leq\mathtt{sre}(w)+2$.
Proof
The lower bound follows from Lemma 1. For the upper bound, we analyse the new right-extensions that may arise due to appending $c$ to $w$. For any fixed suffix $xc$ of $wc$:
  1. 1.
    if $xc$ appears in $w$, or if $xa$ does not appear in $w$ for any $a\neq c$, or both, then $xc$ induces no new right-extensions in $wc$;
  2. 2.
    if for some $a\neq b$, $xa$ and $xb$ were both substrings of $w$, and $xc$ was not, then $xc$ is a new right-extension of $wc$;
  3. 3.
    if $x$ is always followed by $a\neq c$ in $w$ (hence, $xa$ is not a right-extension of $w$), then both $xa$ and $xc$ are new right-extensions of $wc$.
Cases 1 and 2 induce at most one new super-maximal right-extension in total for all possible $xc$, namely the longest right-extension in $wc$ that is a suffix of $wc$. For Case 3, consider a fixed $a\in\Sigma$. For all the increasing-length suffixes $x_{1}c,x_{2}c,\dots,x_{t}c$ of $wc$ that became right-extensions together with $x_{1}a,x_{2}a,\dots,x_{t}a$, one can see that the latter form a chain of suffixes of $x_{t}a$. Hence, we only have one possible new super-maximal right-extension ending with $a$, namely $x_{t}a$. Observe that the chain of suffixes $x_{1}a,x_{2}a,\dots,x_{t}a$ is unique: if the suffix $x$ is always followed by $a$, any suffix $y$ of $x$ is either right-maximal in $w$ (and $y$ falls within Case 2), or it is always followed by an $a$ (because $x$ is always followed by an $a$), i.e. $y=x_{i}$ for some $i\in[1\mathinner{.\,.}t]$. ∎
Lemma 3
Let $w\in\Sigma^{*}$ and $c\in\Sigma$. It holds $\mathtt{sre}(w)\leq\mathtt{sre}(cw)\leq\mathtt{sre}(w)+2$.
Proof
The lower bound follows from Lemma 1. For the upper bound, let cxacxa be the smallest prefix of $cw$ that is not a right-extension of $w$, but is a right-extension of $cw$ (if it exists). This implies that there exists $b\neq a$ such that $cxb\in\mathcal{F}_{w}$ and $cxa\not\in\mathcal{F}_{w}$ (otherwise, cxacxa would be a right-extension of $w)$, so no prefix of $cw$ of length $|cxa|$ or more is right-maximal, and thus no prefix longer than cxacxa can be a right-extension. By the minimality of cxacxa, all prefixes of $cw$ shorter than cxacxa either are not right-extensions of $cw$, or are right-extensions of both $w$ and $cw$. Therefore, cxacxa together with some cxbcxb appearing in $w$, are the only possible new right-extensions in $cw$ with respect to ww.∎
Lemmas 2 and 3 yield a remarkably simple algorithms to compute smallest suffixient sets, which we detail in Section 7.
By letting $c=\text{\tt\textdollar}\not\in\mathcal{F}_{w}$ in Lemma 2, we relate $\chi$ to $\mathtt{sre}$ (note that $\chi$ is always at least $\mathtt{sre}+1$ because of the new super-maximal extension ending with \$). This makes clear the relation between Combinatorics on words [21] with suffixient sets, via the common notion of right-special factors (which we call here right-maximal substrings).
Corollary 1
Let $w\in\Sigma^{*}$. It holds $\mathtt{sre}(w)+1\leq\chi(w)\leq\mathtt{sre}(w)+2$.
Note that, while the value $\mathtt{sre}(w)$ is non-decreasing after appending a character, this is not the case for the measure $\chi$.
Lemma 4
The measure $\chi$ is not monotone to appending a character.
Proof
Let $w=\texttt{abaab}$. It holds $\mathcal{S}_{r}(w\text{\tt\textdollar})=\{\texttt{aa},\texttt{ab},\texttt{ab\text{\tt\textdollar}},\texttt{aba}\}$ and $\mathcal{S}_{r}(w\texttt{a}\text{\tt\textdollar})=\mathcal{S}_{r}(\texttt{abaaba\text{\tt\textdollar}})=\{\texttt{ab},\texttt{aba\text{\tt\textdollar}},\texttt{abaa}\}$. Hence, $\chi(w)=4$ and $\chi(w\texttt{a})=3$. ∎
4.2 Arbitrary edit operations and rotations
Now we study how much $\mathtt{sre}(w)$ can vary upon edit operations at arbitrary positions and rotations. We will use the following famous string family.
Definition 6
A binary de Bruijn sequence of order $k>0$ [2] contains every binary string in $\{\texttt{a},\texttt{b}\}^{k}$ as a substring exactly once. The length of this sequence is $n=2^{k}+(k-1)$. The set of binary de Bruijn sequences of order $k$ is $\mathtt{dB}(k)$.
Lemma 5
It holds $\mathtt{sre}(w)=2^{k}=\Omega(n)$ for any $w[1\mathinner{.\,.}n]\in\mathtt{dB}(k)$.
Proof
Let $w[1\mathinner{.\,.}n]$ be a binary de Bruijn string of order $k$. By definition, $w$ contains every binary string of length $k$ as a substring exactly once. As all the possible pairs of strings $x\texttt{a}$ and $x\texttt{b}$ of length $k$ appear in $w$, it follows that all the strings in $\mathcal{F}_{w}(k)$ are right-extensions. Moreover, each $x\texttt{a}$ and $x\texttt{b}$ of length $k$ are super-maximal right-extensions: otherwise, there would exist some $c\in\{\texttt{a},\texttt{b}\}$ such that $cx\texttt{a}$ and $cx\texttt{b}$ are both substrings of $w$, which raises a contradiction since the $k$-length string $cx$ cannot appear twice in $w$. Moreover, there are no right-maximal strings of length $k$ or greater; hence, there are no right-extensions of length greater than $k$. It follows that $\mathtt{sre}(w)=|\mathcal{F}_{w}(k)|=2^{k}=\Omega(n)$. ∎
The following lemma uses the de Bruijn family to show that $\mathtt{sre}$ can grow by $\Omega(\log n)$ upon arbitrary edit operations and rotations.
Lemma 6
Let $w=\texttt{a}^{k}\texttt{b}\texttt{a}^{k-2}\texttt{b}x\texttt{a}\texttt{b}^{k}\texttt{a}^{k-1}\in\mathtt{dB}(k)$ be the lexicographically smallest binary de Bruijn sequence of order $k$ [11, 12]. It holds:
  1. 1.
    (Ins) $\mathtt{sre}(w)-\mathtt{sre}(w^{\prime})=2k-2$ if $w^{\prime}=\texttt{a}^{2k-2}\texttt{b}x\texttt{a}\texttt{b}^{k}\texttt{a}^{k-1}$,
  2. 2.
    (Sub) $\mathtt{sre}(w)-\mathtt{sre}(w^{\prime})=2k-3$ if $w^{\prime}=\texttt{a}^{k}\texttt{b}\texttt{a}^{k-2}\texttt{b}x\texttt{a}\texttt{b}^{k-1}\texttt{c}\texttt{a}^{k-1}$,
  3. 3.
    (Del) $\mathtt{sre}(w)-\mathtt{sre}(w^{\prime})=2k-4$ if $w^{\prime}=\texttt{a}^{k}\texttt{b}\texttt{a}^{k-2}\texttt{b}x\texttt{a}\texttt{b}^{k}\texttt{c}\texttt{a}^{k-1}$,
  4. 4.
    (Rot) $\mathtt{sre}(w)-\mathtt{sre}(w^{\prime})=2k-2$ if $w^{\prime}=\texttt{b}\texttt{a}^{k-2}\texttt{b}x\texttt{a}\texttt{b}^{k}\texttt{a}^{2k-1}$.
Proof
Observe that in each claim, $w$ is obtained after performing a string operation on the corresponding $w^{\prime}$: in Claim 1, $w\in{\tt ins}_{\Sigma}(w^{\prime})$; in Claim 2, $w\in{\tt sub}_{\Sigma}(w^{\prime})$; in Claim 3, $w\in{\tt del}_{\Sigma}(w^{\prime})$; in Claim 4, $w\in\mathcal{R}(w^{\prime})$. We prove each claim separately by comparing the super-maximal extensions of $w^{\prime}$ before and after performing the string operation on $w^{\prime}$ that yields $w$, for which $\mathtt{sre}(w)=2^{k}$ by Lemma 5.
For Claim 1, note that $\mathtt{sre}(w^{\prime})$ is the same as $\mathtt{sre}(\texttt{a}^{k}\texttt{b}x\texttt{a}\texttt{b}^{k}\texttt{a}^{k-1})$, as prepending the character a multiple times to this string to obtain $w^{\prime}$ never increases $\mathtt{sre}$; it only updates the super-maximal extension $\texttt{a}^{k}$ to $\texttt{a}^{k+1}$ and $\texttt{a}^{k-1}\texttt{b}$ to $\texttt{a}^{k}\texttt{b}$, and so on. For simplicity, we let $w^{\prime}=\texttt{a}^{k}\texttt{b}x\texttt{a}\texttt{b}^{k}\texttt{a}^{k-1}$. The string $w^{\prime}$ does not contain substrings of length $k$ of the form $\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}$ for $i\in[1\mathinner{.\,.}k-2]$, nor the substring $\texttt{b}\texttt{a}^{k-2}\texttt{b}$. Note that for each of these substrings $y\in\mathcal{F}_{w}(k)$ with $y\not\in\mathcal{F}_{w^{\prime}}(k)$, the other corresponding right-extension $y^{\prime}$ in $w$ sharing a length k1k-1 prefix with $y$ is not a right-extension in $w^{\prime}$. Moreover, note that all the suffixes of length k1k-1 of these $y$ are not suffixes of one another, nor of the length k1k-1 suffixes of any of the substrings $y^{\prime}$ in $w^{\prime}$. Hence, all k1k-1 length binary strings still appear in $w^{\prime}$ as the suffix of some length $k$ substring that remains a right-extension in $w^{\prime}$, and hence, super-maximal extensions of $w^{\prime}$ have to be of length at least $k$. As each string of length $k$ appearing in $w^{\prime}$ is unique, there are no super-maximal extensions of length greater than $k$. Thus, $\mathtt{sre}(w^{\prime})=2^{k}-2(k-1)$ because we are losing k1k-1 pairs of super-maximal extensions of length $k$ with respect to $w$. It follows that by inserting the b in $w^{\prime}$ to yield $w$, $\mathtt{sre}$ increases by 2k22k-2.
For Claim 2, note that exactly $k$ substrings of length $k$ are lost when substituting the last b of $w$ by c: those of the form $\texttt{b}^{i}\texttt{a}^{k-i}$ with $i>0$. This means that substrings ending in $\texttt{b}^{i}\texttt{a}^{k-i-1}$ with $0<i<k$ are not right-maximal in $w^{\prime}$, hence, 2(k1)2(k-1) super-maximal extensions are lost. Moreover, $\texttt{b}^{k-2}$ is still a right-maximal substring, since $\texttt{b}^{k-1}$ and $\texttt{b}^{k-2}\texttt{c}$ occur in $w^{\prime}$. Observe that only $\texttt{b}^{k-2}\texttt{c}$ is a super-maximal extension, while $\texttt{b}^{k-1}$ is a suffix of $\texttt{a}\texttt{b}^{k-1}$. Thus, $\mathtt{sre}(w^{\prime})=2^{k}-2(k-1)+1$ and $\mathtt{sre}(w)-\mathtt{sre}(w^{\prime})=2k-3$.
For Claim 3, the analysis is similar to Claim 2, but in $w^{\prime}$, $\texttt{b}^{k-1}$ remains as a super-maximal extension. Thus, $\mathtt{sre}(w^{\prime})=2^{k}-2(k-1)+2$ and $\mathtt{sre}(w)-\mathtt{sre}(w^{\prime})=2k-4$.
For Claim 4, the analysis is similar to Claim 1, but in $w^{\prime}$, $\texttt{b}\texttt{a}^{k-2}\texttt{b}$ appears, while $\texttt{a}^{k-1}\texttt{b}$ does not. Thus, $\mathtt{sre}(w^{\prime})=2^{k}-2(k-1)$ and $\mathtt{sre}(w)-\mathtt{sre}(w^{\prime})=2k-2$. ∎
4.3 Reversals
We now show that $\mathtt{sre}$ can grow by $\Omega(\sqrt{n})$ upon string reversals.
Lemma 7
Let $k>0$. Let $w_{k}=\prod_{i=0}^{k-1}\texttt{c}\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\text{\tt\#}_{i}\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\text{\tt\textdollar}_{i}$ on the alphabet $\Sigma=\{\texttt{a},\texttt{b},\texttt{c}\}\cup\bigcup_{i\in[0\mathinner{.\,.}k-1]}\{\text{\tt\#}_{i},\text{\tt\textdollar}_{i}\}$. It holds $\mathtt{sre}(w_{k})-\mathtt{sre}(w_{k}^{R})=k-1$.
Proof
Observe that by construction, any substring of $w_{k}$ containing $\text{\tt\#}_{i}$ or $\text{\tt\textdollar}_{i}$ is not right-maximal, as these symbols are unique. Hence, the right-extensions of $w_{k}$ cannot cross from one side to the other side of those special delimiters. Moreover, substrings of the form $\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}$ for $i\in[0\mathinner{.\,.}k-1]$ appear exactly twice in $w_{k}$ and their right-extensions are super-maximal. By looking at the structure of the string $w_{k}$ and carefully analyzing its right-extensions, one can verify that the super-maximal right-extensions of $w_{k}$ are the following:
  1. 1.
    $\texttt{b}\texttt{a}^{k-1}$ and c
  2. 2.
    $\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\text{\tt\#}_{i}$ and $\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\text{\tt\textdollar}_{i}$ for $i\in[0\mathinner{.\,.}k-1]$,
  3. 3.
    $\texttt{c}\texttt{a}^{i}$ and $\texttt{c}\texttt{a}^{i-1}\texttt{b}$ for $i\in[1\mathinner{.\,.}k-1]$,
  4. 4.
    $\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}$ for $i\in[1\mathinner{.\,.}k-1]$.
This sums to a total of 5k15k-1 super-maximal extensions in $w_{k}$. In the reversed string $w_{k}^{R}=\prod_{i=0}^{k-1}\text{\tt\textdollar}_{k-i-1}\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\text{\tt\#}_{k-i-1}\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\texttt{c}$, we have instead:
  1. 1.
    $\texttt{b}\texttt{a}^{k-1}$ and $\text{\tt\textdollar}_{k-1}$,
  2. 2.
    $\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\text{\tt\#}_{k-i-1}$ and $\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}\texttt{c}$ for $i\in[0\mathinner{.\,.}k-1]$,
  3. 3.
    $\texttt{a}^{k-i-1}\texttt{c}\text{\tt\textdollar}_{k-i-2}$ for $i\in[1\mathinner{.\,.}k-2]$, and $\texttt{a}^{k-2}\texttt{c}\text{\tt\textdollar}_{k-2}$,
  4. 4.
    $\texttt{a}^{i}\texttt{b}\texttt{a}^{k-i-1}$ for $i\in[1\mathinner{.\,.}k-1]$.
This sums to a total of $4k$ super-maximal extensions in $w_{k}^{R}$. Thus, $\mathtt{sre}(w_{k})-\mathtt{sre}(w_{k}^{R})=(5k-1)-4k=k-1$. This is $\Theta(\sqrt{|w_{k}|})$ since $|w_{k}|=2k^{2}+4k$. ∎
As the construction is quite complex, we give an example of the words $w_{k}$ and $w_{k}^{R}$ of Lemma 7, and their super-maximal right-extensions.
Example 1
Let $w_{3}=\texttt{c}\texttt{baa}\text{\tt\#}_{0}\texttt{baa}\text{\tt\textdollar}_{0}\texttt{c}\texttt{aba}\text{\tt\#}_{1}\texttt{aba}\text{\tt\textdollar}_{1}\texttt{c}\texttt{aab}\text{\tt\#}_{2}\texttt{aab}\text{\tt\textdollar}_{2}$. It can be verified that the super-maximal right-extensions of $w_{3}$ are:
  1. 1.
    baa and c;
  2. 2.
    $\texttt{baa}\text{\tt\#}_{0}$ and $\texttt{baa}\text{\tt\textdollar}_{0}$; $\texttt{aba}\text{\tt\#}_{1}$ and $\texttt{aba}\text{\tt\textdollar}_{1}$; $\texttt{aab}\text{\tt\#}_{2}$ and $\texttt{aab}\text{\tt\textdollar}_{2}$;
  3. 3.
    ca and cb; caa and cab;
  4. 4.
    aba and aab.
Similarly, let $w_{3}^{R}=\text{\tt\textdollar}_{2}\texttt{baa}\text{\tt\#}_{2}\texttt{baa}\texttt{c}\text{\tt\textdollar}_{1}\texttt{aba}\text{\tt\#}_{1}\texttt{aba}\texttt{c}\text{\tt\textdollar}_{0}\texttt{aab}\text{\tt\#}_{0}\texttt{aab}\texttt{c}$. The super-maximal right-extensions of $w_{3}^{R}$ are:
  1. 1.
    baa and $\text{\tt\textdollar}_{2}$;
  2. 2.
    $\texttt{baa}\text{\tt\#}_{2}$ and baac; $\texttt{aba}\text{\tt\#}_{1}$ and abac; $\texttt{aab}\text{\tt\#}_{0}$ and aabc;
  3. 3.
    $\texttt{ac}\text{\tt\textdollar}_{0}$; $\texttt{ac}\text{\tt\textdollar}_{1}$;
  4. 4.
    aba and aab.
One can see that $\mathtt{sre}(w_{3})=14$, $\mathtt{sre}(w_{3}^{R})=12$, and hence, $\mathtt{sre}(w_{3})-\mathtt{sre}(w_{3}^{R})=2$, as stated in Lemma 7.
4.4 Lower and upper bounds on the sensitivity of $\chi$
Formally, the additive sensitivity of a measure of repetitiveness $\mu$ to a string operation $\rho$ can be defined as a function $AS_{\mu,\rho}:\mathbb{Z}^{+}\rightarrow\mathbb{R}$, where $AS_{\mu,\rho}(n)=\max_{w\in\Sigma^{n}}(\max_{w^{\prime}\in\rho(w)}(\mu(w^{\prime}))-\mu(w))$, that is, the maximum achievable difference among all the strings. Overall, we obtain the following result on the additive sensitivity of $\mathtt{sre}$, which, by Corollary 1, can be written in terms of $\chi$.
Corollary 2
The following bounds on the additive sensitivity of the measure $\chi$ to string operations hold:
  1. 1.
    $AS_{\chi,\rho}(n)=\Theta(1)$ for $\rho\in\{\mathtt{append},\mathtt{prepend}\}$;
  2. 2.
    $AS_{\chi,\rho}(n)=\Omega(\log n)$ for $\rho\in\{\mathtt{ins},\mathtt{del},\mathtt{sub},\mathcal{R}(\cdot)\}$;
  3. 3.
    $AS_{\chi,\mathtt{rev}}(n)=\Omega(\sqrt{n})$, where $\mathtt{rev}(w)=\{w^{R}\}$.
Proof
Claim 1 follows by Lemma 6, where $n=|w|=\Theta(2^{k})$ and $AS_{\chi,\rho}(n)=\Omega(k)=\Omega(\log n)$, for all $\rho\in\{\mathtt{ins},\mathtt{del},\mathtt{sub},\mathcal{R}(\cdot)\}$. Claim 2 follows by Lemma 7, where $n=|w_{k}|=\Theta(k^{2})$ and $AS_{\chi,\mathtt{rev}}(n)=\Omega(k)=\Omega(\sqrt{n})$. ∎
Finally, we give upper bounds on the sensitivity of $\chi$ to string operations in terms of measure $\delta$.
Lemma 8
Let $w\in\Sigma^{*}$ and $w^{\prime}\in\mathtt{ins}_{\Sigma}(w)\cup\mathtt{del}_{\Sigma}(w)\cup\mathtt{sub}_{\Sigma}(w)\cup\mathcal{R}(w)\cup\{w^{R}\}$. It holds
$\chi(w^{\prime})-\chi(w)=O\left(\delta\max\left(1,\log(n/(\delta\log\delta))\right)\log\delta\right)\text{ and }$
$\chi(w^{\prime})\hskip 3.0pt/\hskip 4.0pt\chi(w)=O\left(\max\left(1,\log(n/(\delta\log\delta))\right)\log\delta\right).$
Proof
It holds $\delta\leq\chi\leq 2\overline{r}$ [4] and $r=O(\delta\max(1,\log(n/(\delta\log\delta)))\log\delta)$ [17]. Since the multiplicative sensitivity of $\delta$ to any of the string operations is $O(1)$ [1], for any $w\in\Sigma^{*}$ it holds $\overline{r}(w)=r(w^{R})=O(\delta\max(1,\log(n/(\delta\log\delta)))\log\delta)$. The thesis follows by considering the worst case, that is, $\chi(w)=\Theta(\delta)$ and $\chi(w^{\prime})=\Theta(\delta\max(1,\log(n/(\delta\log\delta)))\log\delta)$.∎
5 On the Multiplicative Sensitivity of $\mathtt{sre}$ to Reverse
We prove in this section that $\chi$ at most doubles when we reverse the string.
In the following, we denote the set of symbols preceding a substring $x$ (or left-extensions of $x$) of $w$ by $L_{x}(w)$, and we use $l_{x}(w)$ to denote an arbitrary element of $L_{x}(w)$ (we drop $w$ if it is clear from the context). For simplicity of analysis, we assume that $w$ contains a special symbol \$ at the beginning and at the end. This does not matter in asymptotic terms, as $\mathtt{sre}(w)=\mathtt{sre}(\text{\tt\textdollar}w\text{\tt\textdollar})\pm O(1)$, and similarly, $\mathtt{sre}(\text{\tt\textdollar}w^{R}\text{\tt\textdollar})=\mathtt{sre}(w^{R})\pm O(1)$. We do this to (i) ensure all right-maximal substrings of $w$ have at least two left-extensions, and hence, right-extensions have at least one left-extension, and (ii) to provide the necessary symmetry to prove Lemma 10 below.
Please note that left-extensions are not analogous to right-extensions. A right-extension is a string $xa$ such that $xb$ also occurs for some $b\neq a$, while a left-extension of $x$ is just any symbol preceding $x$. A left-extension can be unique, whereas right-extensions come (at least) in pairs. We will rely on the following observation.
Observation 5.1
If $xa$ is a super-maximal right-extension, then $L_{xa}\cap L_{xb}=\emptyset$ for any $b\neq a$.
Proof
Otherwise, $xa$ is not super-maximal, because there exists $c\in L_{xa}\cap L_{xb}$, for which $cxa\in\mathcal{F}_{w}$ and $cxb\in\mathcal{F}_{w}$, and thus $xa$ is a proper suffix of the right-extension cxacxa of the right-maximal string $cx$. ∎
Let us introduce a special tree of super-maximal extensions, which will be helpful to prove our result.
Definition 7
The prefix super-maximal right-extension tree (PSR tree) of $w$ is a compacted trie of all the super-maximal right-extensions of $w$. Since some of those strings can be prefixes of others, internal nodes of the trie may also represent strings, and their corresponding nodes are not compacted. Distinct leaves do not represent prefixes of each other.
The following result is key for our proof.
Lemma 9
Let $w$ be a string starting with a unique symbol \$. If $x$ is a right-maximal substring of $w$, and $xa$ is a super-maximal right-extension of $x$ in $w$, then $x^{R}$ is a right-maximal substring of $w^{R}$.
Proof
Because $xa$ is a right-extension in $w$, there exists $xb\in\mathcal{F}_{w}$ with $b\neq a$. Any symbol preceding $xa$ cannot precede $xb$, otherwise, $xa$ would not be super-maximal. Since $w$ starts with the unique symbol \$, if $x\neq\varepsilon$, then $xa$ and $xb$ have occurrences that are not prefixes of $w$, and hence, they are preceded by different symbols. It follows $x^{R}$ is right-maximal in $w^{R}$.∎
Lemma 10
It always holds that $\mathtt{sre}(w^{R})/2\leq\mathtt{sre}(w)\leq 2\cdot\mathtt{sre}(w^{R})$.
Proof
We define an injection $\lambda$ from the set of leaves and unary internal nodes (those with exactly one child) of the PSR tree $T(w)$ of $w$ into $\mathcal{S}_{r}(w^{R})$. We will identify tree nodes with the string they represent.
  • Leaf nodes: If $xa$ is (the supermaximal right-extension represented by) a leaf node, we let $\lambda(xa)=ux^{R}\cdot l_{xa}$ where $l_{xa}\in L_{xa}(w)$ (which exists because $w$ starts with \$, as discussed), and $u$ is chosen such that $ux^{R}\cdot l_{xa}\in\mathcal{S}_{r}(w^{R})$. At least one such $u$ must exist because $x^{R}\cdot l_{xa}\in\mathcal{F}_{w^{R}}$ and $x^{R}$ is right-maximal in $w^{R}$ by Lemma 9.
  • Unary internal nodes: If $xa$ a unary internal node with child $yb=xazb$, consider a right-extension xazcxazc of xazxaz, where $c\neq b$ (one must exist because xazxaz is right-maximal). We then let $\lambda(xa)=uz^{R}ax^{R}\cdot l_{yc}\in\mathcal{S}_{r}(w^{R})$ where $l_{yc}\in L_{yc}(w)\subseteq L_{xa}(w)$. Note $L_{yc}$ is not empty because, as in the previous point, xazcxazc is a supermaximal right-extension, and $u$ exists because $z^{R}ax^{R}$ is right-maximal in $w^{R}$ by Lemma 9.
We now prove that $\lambda$ is an injection. Consider two leaves $xa$ and $yb$ of $T(w)$:
  1. 1.
    If $x[i]\neq y[i]$ for some $i$, then $\lambda(xa)=ux^{R}\cdot l_{xa}$ and $\lambda(yb)=vy^{R}\cdot l_{yb}$, aligned from the right, differ at their position $\lambda(xa)[|\lambda(xa)|-i]\neq\lambda(yb)[|\lambda(yb)|-i]$, and therefore are different.
  2. 2.
    If $x$ is a proper prefix of $y$, then and $y$ has to start with $xc$ for some $c\neq a$ (otherwise $xa$ would be internal). Because $xa$ is a supermaximal right-extension, the set of left-extensions $L_{xa}$ and $L_{yb}\subseteq L_{xc}$ are disjoint by Observation 5.1. Therefore, the selected super-maximal extensions are of the form $\lambda(xa)=ux^{R}\cdot l_{xa}$ and $\lambda(yb)=vy^{R}\cdot l_{xc}$ with $l_{xa}\neq l_{xc}$, thus, distinct.
  3. 3.
    Similarly, if $x=y$, the two super-maximal extensions $xa$ and $xb$ have disjoint sets of left-extensions $L_{xa}$ and $L_{xb}$, and hence $\lambda(xa)\neq\lambda(xb)$ because they end in different characters.
For a graphic depiction of these three cases see Figure 2.
In the other cases, one of the two nodes is a (unary) internal node. Assume, without loss of generality, that $|x|\leq|y|$. By the same argument of point 1 above, $\lambda(xa)$ cannot be equal to some $\lambda(yb)$ if $x$ is not a prefix of $y$. It follows that $xa$ is a unary internal PSR node that is an ancestor of $yb=xazb$:
  1. 4.
    Let $yb$ be the child of $xa$. Let $c\neq b$ be such that xazcxazc is a right-extension in $w$, and $\lambda(xa)=uz^{R}ax^{R}\cdot l_{yc}$ with $l_{yc}\in L_{yc}$. Note that $\lambda(xa)$ differs from $\lambda(yb)$ because $\lambda(yb)$ finishes with some $l_{yb}\in L_{yb}$ and, by Observation 5.1, $L_{yb}\cap L_{yc}=\emptyset$.
  2. 5.
    If $yb$ is a farther descendant of $xa$, then xazxaz has children by $c$ and by some $d\neq c$ where xazdxazd prefixes $y$. The string xazdxazd is a supermaximal right-extension because there is a node for it in the tree, hence, by Observation 5.1, it holds $L_{xazc}\cap L_{xazd}=\emptyset$. As $L_{yb}\subseteq L_{xazd}$, it follows $L_{xazc}\cap L_{yb}=\emptyset$. Therefore, $\lambda(xa)$ and $\lambda(yb)$ end in different symbols.
For a graphic depiction of these two cases see Figure 3.
We are then injecting $k$ nodes into $\mathcal{S}_{r}(w^{R})$, where $k\geq\mathtt{sre}(w)/2$ (with equality when $T$ is a full binary tree). Thus, $\mathtt{sre}(w)/2\leq\mathtt{sre}(w^{R})$. Since the reversal operation is an involution, we have similarly that $\mathtt{sre}(w^{R})/2\leq\mathtt{sre}(w)$. ∎
[…]
Figure 2: Cases 1), 2), and 3) of Lemma 10, where $xa$ and $yb$ are leaves of the PSR tree. Note that the suffix $x^{R}\cdot l_{xa}$ and $x^{R}\cdot l_{yb}$ of $\lambda(xa)$ and $\lambda(yb)$ are not a suffix one of the other because the grayed symbols are different.
[…]
Figure 3: Case 4) and 5) of Lemma 10, where $xa$ is an internal node of the PSR tree, ybyb is the only child of $xa$, and $y^{\prime}d$ is a farther descendant of $xa$ (if it exists). Note xazcxazc is a right-extension, but not super-maximal (otherwise, $xa$ would not be unary). In this case, $\lambda(xa)$ is suffixed by $y^{R}\cdot l_{yc}$, whereas $\lambda(yb)$ and $\lambda(y^{\prime}d)$ are suffixed by $y^{R}\cdot l_{yb}$ and $y^{R}\cdot l_{y^{\prime}d}$, respectively. Note how $yb$ and all its descendants yield $\lambda(\cdot)$ values differing from $\lambda(xa)$ in their last symbol.
6 Relating $\chi$ to Other Repetitiveness Measures
Previous work [4] established that $\gamma=O(\chi)$ and $\chi=O(\overline{r})$ on every string family. In this section we obtain the more natural result that $\chi$ is always $O(r)$, and that it can be asymptotically strictly smaller, $\chi=o(r)$, on some string families (we actually prove $\chi=o(v)$). We also show that $\chi$ is incomparable with all the copy-paste measures except $b$, in the sense that there are string families where $\chi$ is asymptotically strictly smaller than each other, and vice versa.
6.1 Proving $\chi=O(r)$
We first prove that $\chi$ is asymptotically upper-bounded by the number $r$ of runs in the BWT of the sequence. As for the measure $\chi$, we assume that the BWT is computed after appending the \$ symbol.
Lemma 11
It always holds that $\chi\leq 2r$.
Proof
Let $x_{i}$ denotes the $i$th rotation of $w\text{\tt\textdollar}$ in lexicographic order, for each $i\in[1\mathinner{.\,.}|w|+1]$, and let $u_{i}$ be the longest common prefix between the rotations $x_{i},x_{i+1}$, for each $i\in[1\mathinner{.\,.}|w|]$ (note this implies that $u_{i}$ is right-maximal). We further define $s:[1\mathinner{.\,.}n+1]\rightarrow[0\mathinner{.\,.}n]$ as $s(i)=j$ if $x_{i}=w[j+1\mathinner{.\,.}|w|]\text{\tt\textdollar}w[1\mathinner{.\,.}j]$, that is, the number of cyclic shifts to the right required to transform $x_{i}$ into $w\text{\tt\textdollar}$. [^2] The function $s$ mimics the well-known Suffix Array [23], here omitted for simplicity of exposition. As the symbol \$ occurs only once in $w\text{\tt\textdollar}$, the function $s$ is bijective.
Note that each right-extension of $w\text{\tt\textdollar}$ can be written as $u_{i}c$, for some $i\in[1\mathinner{.\,.}|w|]$ and $c\in\Sigma$. Consider now the set
$S=\bigcup_{i\in[1\mathinner{.\,.}|w|]}\{s(i)+|u_{i}|+1,s(i+1)+|u_{i}|+1\},$
that is, the set of positions where the occurrences of the right-extensions $u_{i}c_{1}$ and $u_{i}c_{2}$ end in $w\text{\tt\textdollar}$, where $u_{i}c_{1}$ and $u_{i}c_{2}$ are the prefixes of $x_{i}$ and $x_{i+1}$, respectively, for some $c_{1},c_{2}\in\Sigma$ such that $c_{1}<c_{2}$. It follows by construction that the set $S$ is a suffixient set of $w\text{\tt\textdollar}$.
We now show that $|S|\leq 2r$. Let us factorize each pair of consecutive rotations in the BWT-matrix as $x_{i}=u_{i}v_{i}c_{i}$ and $x_{i+1}=u_{i}v^{\prime}_{i}c_{i+1}$. Observe that $v_{i},v^{\prime}_{i}\neq\varepsilon$ [9, Corollary 8], $v_{i}[1]\neq v^{\prime}_{i}[1]$, and $c_{i}=\mathtt{BWT}(w\text{\tt\textdollar})[i]$ for all $i\in[1\mathinner{.\,.}|w|+1]$. A well-known property of the BWT-matrix is that if $c_{i}=c_{i+1}=c\in\Sigma$, then there exists $j\in[1\mathinner{.\,.}|w|]$ such that $x_{j}=cu_{i}v_{i}$ and $x_{j+1}=cu_{i}v^{\prime}_{i}$ [3]. As a consequence, one has that $s(j)+|u_{j}|+1=(s(i)-1)+(|u_{i}|+1)+1=s(i)+|u_{i}|+1$ and $s(j+1)+|u_{j}|+1=(s(i+1)-1)+(|u_{i}|+1)+1=s(i+1)+|u_{i}|+1$, and the procedure can be reiterated as long as $x_{j}$ and $x_{j+1}$ end with the same symbol. It follows that the same set can be written as
$S=\{s(i)+|u_{i}|+1,s(i+1)+|u_{i}|+1\mid i\in[1\mathinner{.\,.}|w|]\wedge\mathtt{BWT}[i]\neq\mathtt{BWT}[i+1]\},$
[…]
6.2 A family with $\chi=o(v)$ (and thus $o(r)$)
We will now show that $\chi=o(v)$ on the so-called Fibonacci words, which also implies $\chi=o(r)$ in that string family because $v=O(r)$ [28]. Combined with Lemma 11, this implies that $\chi$ is a strictly smaller measure than $r$. In contrast, $\chi$ is incomparable with $v$, as we show later. On our way, we obtain some relevant byproducts about the structure of suffixient sets on Fibonacci, and more generally, episturmian words.
Definition 8([8, 16])
An infinite string w is episturmian if it has at most one right-maximal substring of each length and its set of substrings is closed under reversal, that is, $\mathcal{F}_{\textbf{w}}=\mathcal{F}_{\textbf{w}}^{R}$. It is standard episturmian (or epistandard) if, in addition, all the right-maximal substrings of w are of the form $\textbf{w}[1\mathinner{.\,.}i]^{R}$ with $i\geq 0$, that is, they are the reverse of some prefix of w.
Lemma 12
Let $\textbf{w}\in\Sigma^{\omega}$ be an episturmian word with $|\Sigma|=\sigma\geq 2$. Then, $\mathtt{sre}(\textbf{w}[i\mathinner{.\,.}j])\leq\sigma$ for any $i,j\geq 0$.
Proof
Let w be an epistandard word. The right-extensions $x_{1},x_{2},\dots$ ending with $a\in\Sigma$ form a suffix-chain where each $x_{i}$ is a suffix of $x_{i+1}$. This is because the right-maximal substrings are the reverses of the prefixes of the word, so they are all suffixes of the longest one. If we add the same letter to a subset of them to obtain right-extensions, they still form a suffix-chain. There is one of those suffix-chains for each character $a\in\Sigma$.
Let w be episturmian but not necessarily epistandard. There exists some epistandard word s with the same set of substrings, i.e., $\mathcal{F}_{\textbf{w}}=\mathcal{F}_{\textbf{s}}$ [8]. Therefore, for any episturmian word w, there exist exactly $\sigma$ suffix-chains of right-extensions.
When considering substrings of w, the super-maximal right-extension in $\textbf{w}[i\mathinner{.\,.}j]$ ending with $a\in\Sigma$ is the longest right-extension of w ending with $a$ that remains a right-extension in $\textbf{w}[i\mathinner{.\,.}j]$. It follows that for any substring $\textbf{w}[i\mathinner{.\,.}j]$ of any episturmian word w, it holds $\mathtt{sre}(\textbf{w}[i\mathinner{.\,.}j])\leq\sigma$. ∎
Combining this result with Corollary 1, we obtain the following bound.
Corollary 3
For any episturmian word $\textbf{w}\in\Sigma^{\omega}$ it holds $\chi(\textbf{w}[i\mathinner{.\,.}j])\leq\sigma+2$.
The next lemma precisely characterizes the suffixient sets of Fibonacci words, a particular case of epistandard words that will be useful to relate $\chi$ with $v$.
Definition 9
Let $F_{1}=\texttt{b}$, $F_{2}=\texttt{a}$, and $F_{k}=F_{k-1}F_{k-2}$ for $k\geq 3$ be the Fibonacci family of strings. Their lengths, $f_{k}=|F_{k}|$, form the Fibonacci sequence.
Lemma 13
Every Fibonacci word $F_{k}\text{\tt\textdollar}$ has a suffixient set of size at most 4. For $k\geq 6$, the only smallest suffixient sets for $F_{k}\text{\tt\textdollar}$ are $\{f_{k}+1,f_{k}-1,f_{k-1}-1,p\}$, where $p\in\{f_{k-2}+1,2f_{k-2}+1\}$.
Proof
The upper bound of $4$ stems directly from Corollary 3, because the infinite Fibonacci word is binary epistandard. For $k\geq 3$, there exist strings $H_{k}$ such that $F_{k}=F_{k-1}F_{k-2}=H_{k}cd$ and $F_{k-2}F_{k-1}=H_{k}dc$, for $cd=\texttt{a}\texttt{b}$ or $cd=\texttt{b}\texttt{a}$ depending on the parity of $k$ [22]. Let us call $F^{\prime}_{k}=H_{k}dc=F_{k-2}F_{k-1}$, that is, $F_{k}$ with the last two letters exchanged; thus $F_{k}=F_{k-1}F_{k-2}=F_{k-2}F_{k-1}^{\prime}$.
Note that $F_{k-1}=H_{k-1}dc$ prefixes $F_{k}$. On the other hand, we can write $F_{k}=F_{k-1}F_{k-2}=F_{k-2}F_{k-3}F_{k-2}=F_{k-2}F_{k-1}^{\prime}=F_{k-2}H_{k-1}cd$. Therefore, string $H_{k-1}$ is right-maximal in $F_{k}$. Its extensions, $H_{k-1}d$ and $H_{k-1}c$, are super-maximal because there are no other occurrences of $H_{k-1}$ in $F_{k}$: (i) $H_{k-1}$ cannot occur starting at positions $f_{k-2}+2$ or $f_{k-2}+3$ because it occurs at $f_{k-2}+1$, so $H_{k-1}$ should match itself with an offset of 1 or 2, which is impossible because it prefixes $F_{k-1}$ and all $F_{k-1}$ for $k-1\geq 5$ start with abaab; (ii) $H_{k-1}$ cannot occur starting at positions $2$ to $f_{k-2}$ because its prefix $F_{k-2}$ should occur inside the prefix $F_{k-2}F_{k-2}$ of $F_{k}=F_{k-2}F_{k-1}^{\prime}=F_{k-2}F_{k-2}F_{k-3}^{\prime}$, and so $F_{k-2}$ should equal a rotation of itself, which is impossible [7, Cor. 3.2]. The two positions following $H_{k-1}$, $f_{k-1}-1$ and $f_{k}-1$, then appear in any suffixient set.
On the other hand, $F_{k-2}$ is followed by \$ in $F_{k}\text{\tt\textdollar}$, and it also prefixes $F_{k}=F_{k-2}F_{k-1}^{\prime}$, therefore $F_{k-2}$ is right-maximal. The first occurrence is preceded by $F_{k-1}$, and hence by $c$, and the second by no symbol. $F_{k-2}$ also occurs in $F_{k}$ at position $f_{k-2}+1$, as seen above, preceded by $F_{k-2}$ and thus by $d$. There are no other occurrences of $F_{k-2}$ in $F_{k}$ because (i) it cannot occur starting at positions $2$ to $f_{k-2}$ by the same reason as point (ii) of the previous paragraph; (ii) it cannot appear starting at positions $f_{k-2}+2$ to $f_{k-1}-2$ because $F_{k}=F_{k-2}F_{k-2}F_{k-3}^{\prime}$ and $F_{k-3}^{\prime}[1,f_{k-3}-2]=F_{k-3}[1,f_{k-3}-2]=F_{k-2}[1..f_{k-3}-2]$, thus such an occurrence would also match a rotation of $F_{k-2}$, which is impossible as noted above; (iii) it cannot appear starting at positions $f_{k-1}-1$ or $f_{k-1}$ because, since it matches at position $f_{k-1}+1$, $F_{k-2}$ would match itself with an offset of 1 or 2, which is impossible as noted in point (i) of the previous paragraph. The right-extensions of $F_{k-2}$ are then super-maximal. The one followed by \$ occurs ending at position $f_{k}+1$. The other two are followed by a because they are followed by $F_{k-2}$ and by $F_{k-3}^{\prime}$ and all $F_{k}$ for $k\geq 2$ start with a. We can then choose either ending position for a suffixient set, $f_{k-2}+1$ or $2f_{k-2}+1$. ∎
Corollary 4
There exist string families where $\chi=o(v)$.
Proof
It follows from Lemma 13 and the fact that $v=\Omega(\log n)$ on the odd Fibonacci words [28, Thm. 28]. ∎
6.3 Uncomparability of $\chi$ with copy-paste measures
Finally, we show that $\chi$ is incomparable with most copy-paste measures. This follows from $\chi$ being $\Theta(n)$ on de Bruijn sequences and $O(1)$ on Fibonacci strings. Because $g=O(n/\log n)$ on de Bruijn sequences [28] and by Lemma 5, we have:
Corollary 5
There exists a string family with $\chi=\Omega(g\log n)$.
This result is particularly relevant because all the copy-paste based measures $\mu$, with the exception of $z_{e}$, are $O(g)$. Corollary 5 then implies $\mu=o(\chi)$ on de Bruijn sequences for all these measures $\mu$.
While it has been said that $z_{e}=O(n/\log n)$ on binary sequences as well [19], this referred to the version that adds to each phrase the next nonmatching character. Because $z_{e}$ is not an optimal parse, it is not obvious that this also holds for the version studied later in the literature, which does not add the next character. We then prove next that $z_{e}=o(\chi)$ holds on de Bruijn words.
Lemma 14
There exists a string family with $\chi=\Omega\left(z_{e}\frac{\log n\log\log\log n}{(\log\log n)^{2}}\right)$.
Proof
It always holds that $z_{e}=O\left(z\frac{\log^{2}(n/z)}{\log\log(n/z)}\right)$ [13]. In de Bruijn sequences it holds that $z=\Theta(n/\log n)$, so $n/z=\Theta(\log n)$. Therefore, $z_{e}=O\left(z\frac{(\log\log n)^{2}}{\log\log\log n}\right)$, and replacing $z=\Theta(n/\log n)$ we get $z_{e}=O\left(n\frac{(\log\log n)^{2}}{\log n\log\log\log n}\right)$. By Lemma 5, this yields $\chi=\Omega\left(z_{e}\frac{\log n\log\log\log n}{(\log\log n)^{2}}\right)=\omega(z_{e})$ on de Bruijn sequences.∎
Corollary 6
The measure $\chi$ is uncomparable to $\mu\in\{z,z_{no},z_{e},z_{end},v,g,g_{rl},c\}$.
Proof
From Corollary 5 and Lemma 14, and that $z$, $z_{no}$, $z_{end}$, $v$, $g_{rl}$ and $c$ are always $O(g)$, it follows that there are string families where $\mu=o(\chi)$, for any $\mu\in\{z,z_{no},z_{e},z_{end},v,g,g_{rl},c\}$. On the other hand, from Lemma 13 and Corollary 4, and that $c=\Omega(\log n)$ on Fibonacci words [28, Thm. 32] and $c=O(\mu)$ for any $\mu\in\{z,z_{no},z_{e},z_{end},g_{rl},g\}$ [28, Thm. 30], it follows that there are string families where $\chi=o(\mu)$, for any $\mu\in\{z,z_{no},z_{e},z_{end},v,g,g_{rl},c\}$.∎
7 Online Computation of Smallest Suffixient Sets
As an application of Lemmas 2 and 3, we show that Ukkonen’s [33] and Weiner’s [34] linear-time online constructions of suffix trees can be easily modified to compute smallest suffixient sets within the same space and time complexity.
The suffix tree of a text $T$ is a tree of size $O(|T|)$ where edges are labeled by nonempty substrings of $T$ (the label $T[i\mathinner{.\,.}j]$ is indicated by the pair $(i,j)$). Every internal node has at least two children, and the labels of edges to any two children must differ in their first symbol. If node $x$ has a child node $y$ by an edge labeled $(i,j)$, we say that $y$ is a child by label $T[i]$. Each node $x$ is said to be the locus of the string obtained by concatenating the labels of the edges that lead from the root to $x$. The string depth of a node is the length of its locus. The tree leaves are the loci of suffixes $T[i\mathinner{.\,.}]$, whereas internal nodes are loci of strings that occur at least twice in $T$. Edges to leaves have labels of the form $(i,\infty)$, which means the second component is always the last position processed of $T$. Suffix tree nodes also have so-called suffix links, which lead from a node that is the locus of a string $cu$, for $c\in\Sigma$, to the locus of string $u$ (which always exists: if $cu$ occurs more than once, so does $u$). Finally, we extend the concept of suffix tree nodes to implicit nodes, which are virtual nodes with one child, assumed to exist along edges: if $y$ is the child of $x$ by an edge labeled $(i,j)$ and $x$ is the locus of string $u$, then $(x,(i,p))$, for $i\leq p<j$, is an implicit node that is the locus of $u\cdot T[i\mathinner{.\,.}p]$. Knowing the loci of which string they are, suffix links are also defined from implicit nodes (those can lead to explicit or to implicit nodes).
7.1 Ukkonen’s left-to-right construction
Ukkonen’s algorithm [33] builds the suffix tree of $T$ such that, after having processed any prefix $w$ of $T$, it has built the suffix tree of $w$. To prevent special cases, it assumes that the root is in fact a child of a special node $\bot$, with edges to the root labeled $c$ for every $c\in\Sigma$; the suffix link of the root points to $\bot$. We do not aim at a full explanation of the algorithm, but just highlight some of its relevant properties. Let the prefix $w$ of $T$ be followed by $c\in\Sigma$. The algorithm maintains pointers to two (possibly implicit) nodes of the suffix tree of $w$:
$s$:
called the active point of $w$, is the locus of the longest suffix $u$ of $w$ that occurs at least twice in $w$;
$s^{\prime}$:
called the end point of $w$, is the locus of the longest suffix $v$ of $w$ such that $vc$ occurs in $w$.
Note that this implies that the locus of $vc$ is the active point of $wc$ (i.e., the new $s$ after processing $c$), and that, because $v$ must be a suffix of $u$, there is a chain of consecutive suffix links from $s$ to $s^{\prime}$. The algorithm updates the suffix tree nodes in that chain, from $s$ to (but not including) $s^{\prime}$, by adding a new leaf child labeled $c$ to those nodes. Although updating the suffix tree of $w$ to obtain that of $wc$ may take non-constant time, the time does amortize to constant [33].
We will enhance the suffix tree nodes by adding a mark to the nodes that are loci of the supermaximal right-extensions of a smallest suffixient set of $w$. The length of the extension is the string depth of the marked node, and any suffix descending from the node serves as a starting position of such extension. This is the way in which we maintain a smallest suffixient set for the current prefix $w$ of $T$. For example, we can easily maintain the marked nodes in a list in order to collect the set for any prefix in optimal time. To compute $\chi(T)$, we can run the algorithm on $T\text{\tt\textdollar}$ and then return the length of the list.
Note that the loci of right-extensions are either explicit nodes, or implicit nodes of the form $(x,(i,i))$, because they extend by one symbol a string that occurs more than once. We can then always associate the mark to the corresponding explicit child $y$ of $x$, so as to consult and update it in constant time.
The first letter $c$ of $T$ is processed by adding a child leaf of the root by label $c$ and setting $s$ to the root. From there on, considering the cases of Lemma 2, we proceed as follows:
$s$ has a child by label $c$:
This means that $s=s^{\prime}$ and Ukkonen’s algorithm just descends by $c$ to find the new $s$. We do not need any further action because we are in case 1 of the lemma.
$s$ has two or more children, none by label $c$:
This is case 2 of the lemma. Ukkonen’s algorithm will create a new leaf child by label $c$ of $s$, and of all the nodes in the suffix-link chain until it finds $s^{\prime}$ (i.e., the first node in the chain having a child by label $c$). We then (1) mark the (just created) child of $s$ by label $c$, and (2) unmark, if it is marked, the (already existing) child of $s^{\prime}$ by label $c$. This is because $uc$ is a new supermaximal right extension of the right-maximal string $u$ and, in case (2), $vc$ ceases to be a supermaximal right extension because it is a suffix of the new one we are adding, $uc$.
$s$ has just one child, by label $a\neq c$:
This is case 3 of the lemma. Ukkonen’s algorithm proceeds exactly as in the previous case, finding the first $s^{\prime}$ with a child by label $c$ in the suffix-link chain. We (1) mark the two children of $s$ (by labels $a$ and $c$), (2) unmark, if it is marked, the child of $s^{\prime}$ by label $c$, and (3) unmark, if it is marked, the child of $s^{\prime\prime}$ by label $a$, where $s^{\prime\prime}$ is the first node in the suffix-link chain that has another child labeled $b\neq a$ (note that $s^{\prime\prime}$ must be on the chain from $s$ to $s^{\prime}$, because one possible choice is $b=c$ and $s^{\prime\prime}=s^{\prime}$). This is because $ua$ and $uc$ are new supermaximal right-extensions of $u$, and in case (2), $vc$ is not anymore supermaximal, as before. In case (3), similarly, if $s^{\prime\prime}$ is the locus of $z$, then $za$ is not anymore supermaximal. Note that the child by label $a$ of $s^{\prime\prime}$ is the only locus of some $za$ suffix of $ua$ that could possibly be marked.
Ukkonen’s algorithm is claimed to be $O(n)$ time, but this assumes that the alphabet is constant. If it is not, we may need more time to find the child by label $c$ among its children, or determine it does not exist. If the alphabet is an integer range and of size polynomial in $n$, we can still retain $O(n)$ time in the transdichotomous RAM model of computation with computer word size $\omega=\Omega(\log n)$ using fusion trees, as the children operations can then be handled in time $O(\log_{\omega}|\Sigma|)=O(1)$ [31]. Otherwise, an extra factor of $O(\log|\Sigma|)$ appears.
Theorem 7.1
There exists an algorithm to compute smallest suffixient sets that processes a text $T$ left to right such that, for every $n$, after having processed the prefix $T[1\mathinner{.\,.}n]$ it has determined a smallest suffixient set of $T[1\mathinner{.\,.}n]$ using $O(n)$ space and $O(n)$ worst-case time. This can be used to compute $\chi(T)$ within $O(|T|)$ space and time.
7.2 Weiner’s right-to-left construction
Weiner’s algorithm [34] process the text right-to-left, prepending a symbol at each step. Analogously to Ukkonen’s algorithm, after having processed a suffix of $T$, it has built the suffix tree of that suffix. The algorithm is possibly less convenient than Ukkonen’s, as it is more natural to process the text left to right, but its adaptation to compute $\mathtt{sre}$ using Lemma 3 needs diving less into its details. Further, it has the property that the smallest suffixient set it computes for a suffix of $T$ is a subset of those it computes for the following (longer) suffixes, and that we immediately know $\chi$ for each suffix of $T$ as soon as we process it.
[…]
While the algorithm manages to find $v$ in constant amortized time from $u$ using (and maintaining) the so-called “Weiner links” (which are the inverse of the suffix links), let us reason as if we found $v$ by descending from the root with the symbols of $cw\mathdollar$. We would descend by the path until we find an explicit or implicit node $v$ that is the locus of a prefix $cx$ of $cw\mathdollar$ from where no child descends by label $a$, where cxacxa also prefixes $cw\mathdollar$. Those are the strings $cx$ and cxacxa alluded to in Lemma 3. The algorithm then:
  • If $v$ is implicit, converts it into an explicit node with (temporarily) a single child, which descends by label, say, $b$.
  • Creates a new child of $v$ by label $a\neq b$, which is a leaf representing $cw\mathdollar$.
[…]
As Ukkonen’s, Weiner’s algorithm is linear-time under the same assumptions of a transdichotomous RAM model with computer word size $\Omega(\log n)$; otherwise a factor of $O(\log|\Sigma|)$ multiplies the time complexity.
Theorem 7.2
There exists an algorithm to compute smallest suffixient sets that processes a $T\mathdollar$ right to left such that, for every $n$, after having processed the suffix $w[1\mathinner{.\,.}n]\mathdollar$ of $T\mathdollar$ it has determined a smallest suffixient set of $w\mathdollar$, and its size $\chi(w)$, using $O(n)$ space and $O(n)$ worst-case time.
8 Conclusions and Open Questions
We have contributed to the understanding of $\chi$ as a new measure of repetitiveness, better finding its place among more studied ones. Figure 1 shows the (now) known relations around $\chi$ (cf. [27]). As a direct consequence of our findings, we derive new simple online algorithms to compute smallest suffixient sets, and thus $\chi$, as a modification of Ukkonen’s and Weiner’s suffix tree constructions [33, 34].
There are still many interesting open questions about $\chi$. One of the most important is whether $\chi$ is reachable. Proving $b=O(\chi)$ would settle this question on the affirmative, and at the same time give the first copy-paste measure that is comparable with $\chi$. We conjecture, instead, that $\chi$ is not reachable, proving which would imply that $\gamma$ is also unreachable, a long-time open question.
One consequence of Corollary 5 is that $\chi\not\in O(g\log^{k}(n/g))$ for any $k>0$. It could be the case, though, that $\chi\in O(\delta\log n)$, because the separation of $\chi$ and $\delta$ on de Bruijn sequences is a $\Theta(\log n)$ factor.
Regarding edit operations, experiments suggest that $\mathtt{sre}(w^{\prime})/\mathtt{sre}(w)$ is $O(1)$ for all the string operations we considered. We proved such a multiplicative constant for reversals. Proving the same for insertions would imply the existence of a constant for rotation and vice versa. It is also open whether $r=O(\chi\log\chi)$. If this were true —and provided that $\chi$ has $O(1)$ multiplicative sensitivity to string operations— it would imply that $r$ has $O(\log n)$ multiplicative sensitivity to these operations, making the already known lower bounds on multiplicative sensitivity [1, 14, 15] tight. If the conjecture were false, then $\chi$ could be considerably smaller than $r$ in some string families. In the same line, observe that Lemma 10 implies that, either $r(w)/r(w^{R})=\omega(\log n)$ and $r$ can be $\omega(\chi\log n)$, or $r(w)/r(w^{R})=O(\log n)$ and this known bound is tight.
Acknowledgements
We thank Davide Cenzato, Nicola Prezza, and Francisco Olivares for their code to compute smallest suffixient sets https://github.com/regindex/suffixient [5], which was helpful to propose and discard hypotheses on the behavior of $\chi$, for useful discussions on suffixient sets, and to verify the implementation of our Ukkonen’s extension that computes them online.
Disclosure of interests
The authors have no competing interests to declare that are relevant to the content of this article.
Funding
G.N. and C.U. were partially funded by Basal Funds FB0001 and AFB240001, ANID, Chile; and FONDECYT Project 1-230755, ANID, Chile.
G.R. was partially funded by the MUR PRIN Project “PINC, Pangenome INformatiCs: from Theory to Applications” (Grant No. 2022YRB97K), funded by Next Generation EU PNRR M4 C2, Inv. 1.1 and by the INdAM - GNCS Project CUP_\_E53C24001950001.
C.U. was partially funded by ANID-Subdirección de Capital Humano/Doctorado Nacional/2021-21210580, ANID, Chile; and NIC Chile Doctoral Scholarship, NIC, Chile.